# Variational Source-Channel Coding for Semantic Communication

### v2

@@ Abstract
The rapid development of artificial intelligence (AI) brings new opportunities and challenges to classical communication. AI applications not only impose significant transmission pressure on current communication networks but also affect the way information is stored and utilized due to its potent nonlinear fitting capabilities. Consequently, semantic communication technology emerges as a pivotal bridge connecting AI with classical communication. The current semantic communication systems are generally modeled as an Auto-Encoder (AE). AE lacks a deep integration of AI principles with communication strategies due to its inability to effectively capture channel dynamics. This gap makes it difficult to justify the need for joint source-channel coding (JSCC) and to explain why performance improves. This paper begins by exploring lossless and lossy communication, highlighting that the inclusion of data distortion distinguishes semantic communication from classical communication. It breaks the conditions for the separation theorem to hold and explains why the amount of data transferred by semantic communication is less. Therefore, employing JSCC becomes imperative for achieving optimal semantic communication. Moreover, a Variational Source-Channel Coding (VSCC) method is proposed for constructing semantic communication systems based on data distortion theory, integrating variational inference and channel characteristics. Using a deep learning network, we develop a semantic communication system employing the VSCC method and demonstrate its capability for semantic transmission. We also establish semantic communication systems of equivalent complexity employing the AE method and the VAE method. Experimental results reveal that the VSCC model offers superior interpretability compared to AE model, as it clearly captures the semantic features of the transmitted data, represented as the variance of latent variables in our experiments. In addition, VSCC model exhibits superior semantic transmission capabilities compared to VAE model. At the same level of data distortion evaluated by PSNR, VSCC model exhibits stronger human interpretability, which can be partially assessed by SSIM.
@@ I Introduction
Building on these studies, it was recognized that traditional statistical theory-based methods might not fully capture semantic information. Researchers have begun to explore ANNs’ potential to integrate semantic information, with the continuous refinement of AI theory and the maturation of AI technologies. Theoretically, Kolchinsky et al. [24] defined semantic information as “the syntactic information that a physical system has about its environment which is causally necessary for the system to maintain its own existence”. Kountouris et al. [25] envision a communication paradigm shift based on AI, which makes the semantics of information the foundation of the communication process. Their definition entails a goal-oriented unification of information generation, transmission, and reconstruction, by taking into account process dynamics, signal sparsity, data correlation, and semantic information attributes. Yuan et al. [26] also employ AI to study semantic information. Korner entropy is employed to describe the number of AI states and semantic information [27]. Technically, numerous semantic communication systems have been developed based on various ANN models. The study initially appeared with joint source-channel coding (JSCC) based on deep learning, tracing back to Li et al. [28], in 2003. They attempted to replace source coding, channel coding, and modulation with Fully Connected Networks (FCNs). However, due to the limited development of AI technology at that time, the experimental results of the communication system were unsatisfactory. Subsequently, in 2019, Farsad et al. [29] incorporated a communication channel into the Seq2Seq model based on Long Short-Term Memory (LSTM) for JSCC in text transmission, achieving superior performance compared to traditional coding methods. The DeepJSCC system for image transmission was developed in 2020 by Deniz et al. [30], based on Convolutional Neural Networks (CNNs). This model, essentially a DCGAN, was trained using Auto-Encoder (AE) loss functions and exhibited performance comparable to classical communication while avoiding cliff effects. Based on these studies, models such as DeepJSCC-f [31], DeepJSCC-q [32], DeepJSCC-1++ [33] and JSCCformer-f [34] were developed to address feedback and modulation in practical communication. The concept of semantics was first introduced to explain JSCC gains in 2021 by Xie et al. [35]. They built DeepSC model based on Transformer and FCN for text transmission, leading to the emergence of numerous similar models termed semantic communication systems. Semantic communication models have since been developed into various structures targeting different data types, such as the NTSCC (Nonlinear Transform Source Channel Coding) model for image transmission [36] and the SVC (Semantic Video Conferencing) model for video transmission [37].
@@
In response to these three questions, this paper proposes a corresponding theoretical analysis. Based on the analysis, a new JSCC method and loss function are derived using variational inference. An ANN model trained using this method, and the experimental results demonstrated the effectiveness of the approach.
Addressing the first question, drawing from prior research on semantic measurement, this paper posits that data information and semantic information are not one-to-one correspondences. Data serves as the carrier for semantic information, but the entirety of data information is not necessary to convey semantics. For instance, consider the sentences “Bai is not a cat, it’s a dog” and “Bai is a dog”, where the former contains more data information than the latter, yet both convey the same semantics. Therefore, in semantic communication, data distortion at the source is necessary [40]. This contradicts the source coding theorem in CIT, which aims for lossless communication. Consequently, we define semantic communication as allowing data distortion [41][42], yet achieving semantic fidelity through the establishment of prior information (knowledge base). Semantic communication cannot be effectively evaluated with bit error rate [43]. It is based on rate-distortion theorem, and exhibits greater tolerance for data errors, which explains its main differences from classical communication and the gain source.
Building on the investigation of the first question, this paper further explores the second question. While the lossy transmission theory is not yet fully developed, existing proofs indicate that JSCC can still achieve optimality [44]. JSCC considers not only the source and channel noise distributions but also the relationship between source distortion and channel noise. In most cases of source distortion, the source encoding rate and channel encoding rate in separated source-channel coding (SSCC) lack closed-form analytical expressions, making it impossible to achieve optimality in practice [45]. Therefore, JSCC can provide advantages over SSCC when striving for optimal transmission under practical constraints. In semantic communication, the need to account for the knowledge base and the effectiveness problem further enhances the advantages of JSCC.
The exploration of the two preceding questions elucidates that effective semantic communication requires explicit channel modeling in the JSCC. Since semantic communication inherently involves data distortion to eliminate redundant information unnecessary for semantic expression, lossy encoding is applied at the source. According to the optimal theorem for JSCC under lossy conditions, the lossy source in semantic communication could be effectively achieved through the channel noise. Consequently, the channel no longer acts as an obstacle that necessitates additional data redundancy, such as channel encoding, to offset its impact, but instead becomes an integral component of the semantic coding process. To effectively utilize channel for distorting source data, appropriate matching between the channel noise distribution and the source data distribution is necessary. This alignment essentially transforms the original source data distribution into a new distribution, facilitating controlled distortion of data through the noise distribution, while simultaneously preserving the essential characteristics of the original data distribution. Nevertheless, the new distribution may lack a closed-form analytical expression, because data distribution transformation has always been a challenge in statistics [46]. Fortunately, research on ANNs offers effective approximation methods for addressing these challenges [47][48]. Building on variational inference, this paper proposes a Variational Source-Channel Coding (VSCC) method that integrates channel state information into the training function of semantic communication, thereby achieving the necessary alignment process.
@@
The remainder of this paper is organized as follows. In the next section, a detailed explanation of the relationship between semantic communication and classical communication is provided, demonstrating the rationale for implementing semantic communication based on the JSCC. Building upon this research, a VSCC method matched to the channel is proposed. In Section III, the VSCC model implemented using ANNs is elaborately introduced, along with the associated training and testing algorithms. Section IV offers a comprehensive overview of the experimental results based on the proposed VSCC method, in comparison with the AE method and the VAE method. And a careful analysis of the experimental results is conducted, providing explanations and demonstrating the validity of the semantic communication definitions. In Section V, the entire paper is summarized, while also highlighting potential shortcomings in the current experiments and identifying directions for future research.
II Theoretical Analysis
@@
II-A Lossless and Lossy Transmission
Based on previous studies in semantic information theory, there exists a mutual dependency between semantics and data. Semantics invariably depend on the existence of data, but data does not necessarily convey semantics. When the semantic encoder extracts semantic features from data, it removes data redundancies, making semantic communication potentially more efficient than classical communication methods. Therefore, the presence or absence of data distortion is the primary distinction between semantic communication and classical communication. Building on this premise, this paper argues that semantic communication functions within the framework of rate distortion theory for data, achieving semantic fidelity in information transmission through the effective utilization of the knowledge base. This assertion highlights the need to distinguish between lossless communication and lossy communication.
Figure 1: Communication model based on classical communication theory. It mainly consists of a source module, an encoding module, a channel module, a decoding module, and a destination module. The encoding module is divided into source encoding and channel encoding, which are guaranteed by Shannon’s source coding theorem, Shannon’s channel coding theorem, and the source-channel separation theorem.
Classical communication is modeled on the source coding theorem, channel coding theorem, and source-channel coding theorem. As shown in Fig. 1, let $X$ denotes the source, $P(X)$ denotes the source distribution, $H(X)$ denotes the source entropy, and $R$ denotes the coding rate. Assuming that the source symbols belong to the typical set and disregarding their specific meaning, the Shannon’s source coding theorem can be derived.
@@ II-A Lossless and Lossy Transmission
The source coding theorem shows that for a given discrete memoryless source (DMS), if the coding rate $R$ exceeds the source entropy $H(X)$, there exists a coding method make the source decoding error rate arbitrarily small as the code word length approaches infinite. However, practical sources are generally not DMS (especially for images), code word length cannot be infinite, and there are inherent limitations on decoding complexity.
@@
The channel coding theorem indicates that when the encoding rate $R$ is less than the channel capacity $C$, there exists a coding method that can make the channel decoding error rate arbitrarily small as the codeword length approaches infinity. This error rate decreases exponentially with the increase in code length. However, it is generally impossible to ensure that a practical channel is DMC. Channel coding length cannot be infinitely long, and channel decoding cannot have negligible complexity.
It is worth noting that the aforementioned source coding theorem is independent of the channel noise distribution, and the channel coding theorem is independent of the source data distribution. Therefore, according to Shannon’s source-channel coding theorem, the conclusion is reached that if the original source is transmitted without distortion, SSCC is equivalent to JSCC, and both can achieve optimal transmission.
@@
The source-channel coding theorem also relies on the assumptions that the source is memoryless, stationary, and ergodic, and that the channel is DMC, without considering the complexity of encoding and decoding. In addition, these theorems are designed for lossless data transmission. When distortion is introduced to the original source, the communication process can no longer be adequately described by these three theorems alone.
In semantic communication, these assumptions are not fully applicable, particularly the requirement for losslessness. Instead, the semantic coding should obey the rate-distortion theorm.
@@
In practical communication processes, source encoding is inherently lossy. This is due to the impracticality of accurately calculating the actual data entropy, especially for images. The challenge arises because the data being transmitted typically cannot effectively represent the entire population, and thus the estimated distribution does not accurately reflect the population distribution. Moreover, the goal of practical communication is generally to achieve better transmission performance in typical communication scenarios rather than to optimize over an infinitely long time. As a result, data is not considered from the perspective of a typical set, and the source does not exhibit the asymptotic equipartition property (AEP). The effectiveness of data compression and recovery is not simply measured by the bit error rate, but rather by evaluating the source encoding and decoding in terms of the content of the data within the current communication process. Therefore, even without considering channel matching, semantic communication has already been “applied” in practical communication processes.
II-B Joint and Separated Source-Channel Coding
@@
The data transmission described by the rate-distortion function no longer focuses solely on matching the source distribution $P(X)$ with the channel transition distribution $P(\hat{Y}|Y)$, as in CIT. Instead, it integrates the distortion function $d(x,\hat{x})$, source distribution $P(X)$, and channel transition distribution $P(\hat{Y}|Y)$. Consequently, optimal coding must not only minimize the distortion rate but also consider whether the channel noise can be leveraged as part of the distortion function. Subsequent studies have elaborated on this approach.
@@
However, is JSCC truly indispensable for semantic communication? Recent studies have shown that finding an analytical solution for the source distortion rate that perfectly matches a fixed channel coding rate, or vice versa, is unattainable. Unlike CIT, which benefits from the separation theorem, semantic communication lacks a well-established mathematical framework to support such endeavors.
@@
As depicted in Fig.2, we propose a semantic communication model based on JSCC. While semantic information $S$ is transmitted in the data space, its theoretical measurement occurs in the semantic space. The data transmission model resembles the classical communication model, but with SSCC replaced by JSCC, potentially providing a better alignment between source data distortion $d(X;\hat{X})$ and channel noise. Additionally, semantic communication involves the deployment of a knowledge base and the consideration of communication objectives within the semantic space. The knowledge base, serving as a repository of prior information, bolsters the semantic encoding and decoding processes. Furthermore, for diverse communicative tasks, semantic information may also introduce distortions $d_{D}(S;\hat{S})$, retaining only the information relevant to the task, which suggests that semantic communication has the potential to further reduce data transmission. Consequently, achieving optimal semantic communication requires a delicate balance among various factors, including data distortions, source distribution, channel transition distribution, the knowledge base, and semantic distortions. While traditional statistical theories lack the necessary tools to analyze such complex relationships, leveraging JSCC based on ANNs may offer a viable approximation.
II-C Variational Source-Channel Coding for Channel Matching
In a semantic communication system based on ANNs, the joint encoder and decoder are typically trained using the AE method, with the channel functioning as a layer to transmit the latent variables. However, the AE method is not sufficiently precise for communication purposes, as it lacks channel modeling and cannot adequately explain the data distortion process. To address this, we propose a variational source-channel coding (VSCC) method based on the variational inference algorithm. This approach provides a JSCC framework for aligning the data distortion process with the channel.

### v3

@@ Abstract
The rapid development of artificial intelligence (AI) brings new opportunities and challenges to classical communication. AI applications not only impose significant transmission pressure on current communication networks, but also affect the way information is stored and utilized because of its potent nonlinear fitting capabilities. Consequently, semantic communication technology emerges as a pivotal bridge connecting AI with classical communication. The current semantic communication systems are generally modeled as an Auto-Encoder (AE), while it lacks a deep integration of AI principles with communication strategies due to its inability to effectively capture channel dynamics. This gap makes it difficult to justify the need for joint source-channel coding (JSCC) and explain why performance improves. This paper begins by exploring lossless and lossy communication, highlighting that the inclusion of data distortion distinguishes semantic communication from classical communication. It breaks the conditions for the separation theorem to hold and explains why the amount of data transferred by semantic communication is less. Therefore, employing JSCC becomes imperative for achieving optimal semantic communication. Moreover, a Variational Source-Channel Coding (VSCC) method is proposed for constructing semantic communication systems based on data distortion theory, integrating variational inference and channel characteristics. Using a deep learning network, we develop a semantic communication system employing the VSCC method and demonstrate its capability for semantic transmission. We also establish semantic communication systems of equivalent complexity employing the AE method and the Variational ̵Auto-Encoder (VAE) method. The experimental results reveal that the VSCC model offers superior interpretability compared to the AE model, as it clearly captures the semantic features of the transmitted data, represented as the variance of latent variables in our experiments. In addition, the VSCC model exhibits superior semantic transmission capabilities compared to the VAE model. At the same level of data distortion evaluated by PSNR, the VSCC model exhibits stronger human interpretability, which can be partially assessed by SSIM.
@@ I Introduction
Building on these studies, it was recognized that traditional statistical theory-based methods might not fully capture semantic information. Researchers have begun to explore ANNs’ potential to integrate semantic information, with the continuous refinement of AI theory and the maturation of AI technologies. Theoretically, Kolchinsky et al. [24] defined semantic information as “the syntactic information that a physical system has about its environment which is causally necessary for the system to maintain its own existence”. Kountouris et al. [25] envision a communication paradigm shift based on AI, which makes the semantics of information the foundation of the communication process. Their definition entails a goal-oriented unification of information generation, transmission, and reconstruction, by taking into account process dynamics, signal sparsity, data correlation, and semantic information attributes. Yuan et al. [26] also employ AI to study semantic information. Körner entropy is employed to describe the number of AI states and semantic information [27]. Technically, numerous semantic communication systems have been developed based on various ANN models. The study initially appeared with joint source-channel coding (JSCC) based on deep learning, tracing back to Li et al. [28], in 2003. They attempted to replace source coding, channel coding, and modulation with Fully Connected Networks (FCNs). However, due to the limited development of AI technology at that time, the experimental results of the communication system were unsatisfactory. Subsequently, in 2019, Farsad et al. [29] incorporated a communication channel into the Seq2Seq model based on Long Short-Term Memory (LSTM) for JSCC in text transmission, achieving superior performance compared to traditional coding methods. The DeepJSCC system for image transmission was developed in 2020 by Deniz et al. [30], based on Convolutional Neural Networks (CNNs). This model, essentially a DCGAN, was trained using Auto-Encoder (AE) loss functions and exhibited performance comparable to classical communication while avoiding cliff effects. Based on these studies, models such as DeepJSCC-f [31], DeepJSCC-q [32], DeepJSCC-1++ [33] and JSCCformer-f [34] were developed to address feedback and modulation in practical communication. The concept of semantics was first introduced to explain JSCC gains in 2021 by Xie et al. [35]. They built DeepSC model based on Transformer and FCN for text transmission, leading to the emergence of numerous similar models termed semantic communication systems. Semantic communication models have since been developed into various structures targeting different data types, such as the NTSCC (Nonlinear Transform Source Channel Coding) model for image transmission [36] and the SVC (Semantic Video Conferencing) model for video transmission [37].
II Problem Definition and Contributions
II-A Problem Definition
@@ II-B Contributions
In response to these three questions, this paper proposes a corresponding theoretical analysis. Based on the analysis, a new JSCC method and loss function are derived. An ANN model trained using this method, and the experimental results demonstrated its effectiveness. Consequently, this paper makes the following key contributions:
  1. 1
    By formally defining the core principle of tolerable data distortion, this work establishes a theoretical foundation for semantic communication, marking a fundamental departure from classical communication paradigms.
  2. 2
    Through rigorous analysis grounded in CIT, the Lossy Coding Theorem, and the Information Bottleneck (IB) principle, we show that semantic communication must operate within the Joint Source-Channel Coding (JSCC) framework to attain optimal performance. Based on this insight, we further propose an semantic communication model.
  3. 3
    A novel Variational Source-Channel Coding (VSCC) method is proposed that models the channel based on VAEs. In contrast to conventional VAEs, where the encoder deterministically generates latent representations, the proposed method incorporates the physical communication channel as an integral part of the encoding process via IB principle.
  4. 4
    The VSCC model is implemented using ResNet Blocks and attention mechanisms, and is evaluated against AE-based and VAE-based semantic communication models with comparable parameter scales. The AE-based model serves as a representative of conventional semantic communication, while the VAE-based model acts as the algorithmic baseline for VSCC.
  5. 5
    Experimental results show that semantic features remain consistent across varying transmitted data, yet these invariant features necessitate dynamic adaptation to fluctuating channel. It indicates that the latent representations in the VSCC model emerge naturally from channel characteristics, rather than being artificially constructed by the encoder like VAE.
  6. 6
    Experimental results show that the optimal Channel Matching Coefficient (CMC) in the loss function varies with the channel SNRs, suggesting that the channel functions as a part of the joint encoder and contributes to the overall data distortion. The proposed VSCC framework balances the trade-off between semantic distortion and channel noise by dynamically adapting to both the source characteristics and channel conditions.
  7. 7
    Experimental results show that the proposed VSCC model consistently outperforms VAE-based approaches in terms of both Peak Signal-to-Noise Ratio (PSNR), which primarily reflects reconstruction quality, and Structural Similarity Index (SSIM), which serves as a proxy for semantic fidelity. Moreover, under challenging channel conditions, VSCC retains performance advantages over conventional AE-based semantic communication models.
@@
The remainder of this paper is organized as follows. In the next section, the aforementioned three issues will be addressed. A detailed explanation of the relationship between semantic communication and classical communication is provided, demonstrating the rationale for implementing semantic communication based on the JSCC. Building upon this research, a VSCC method matched to the channel is proposed. In Section IV, the VSCC model implemented using ANNs is elaborately introduced, along with the associated training and testing algorithms. Section V offers a comprehensive overview of the experimental results based on the proposed VSCC method, in comparison with the AE method and the VAE method. And a careful analysis of the experimental results is conducted, providing explanations and demonstrating the validity of the semantic communication definitions. In Section VI, the entire paper is summarized, while also highlighting potential shortcomings in the current experiments and identifying directions for future research.
III Theoretical Analysis
@@
III-A Lossless and Lossy Transmission
To address the first research question outlined in Section I, this paper builds upon existing semantic measurement studies to propose a fundamental assumption that data information and semantic information are not one-to-one correspondences. Data serves as the carrier for semantic information, but not the entirety of data information is necessary to convey semantics. For instance, consider the sentences “Bai is not a cat, it’s a dog” and “Bai is a dog”, where the former contains more data information than the latter, yet both convey the same semantics. Therefore, in semantic communication, data distortion at the source is necessary [40]. This contradicts the source coding theorem in CIT, which aims for lossless communication. Semantic communication cannot be effectively evaluated with bit error rate [43]. It is based on rate-distortion theorem, and exhibits greater tolerance for data errors, which explains its main differences from classical communication and the gain source. Consequently, we define semantic communication as allowing data distortion [41][42], yet achieving semantic fidelity through the establishment of prior information (knowledge base). And this assertion highlights the need to distinguish between lossless communication and lossy communication.
Figure 1: Communication model based on classical communication theory. Basically, it consists of a source module, an encoding module, a channel module, a decoding module, and a destination module. The encoding module is divided into source encoding and channel encoding, which are guaranteed by Shannon’s source coding theorem, Shannon’s channel coding theorem, and the source-channel separation theorem.
As shown in Fig. 1, let $X$ denote the source, $P(X)$ denote the distribution of the source, $H(X)$ denote the entropy of the source, and $R$ denote the rate of coding. Assuming that the source symbols belong to the typical set and disregarding their specific meaning, Shannon’s source coding theorem can be derived.
@@ III-A Lossless and Lossy Transmission
It is worth noting that the aforementioned source coding theorem is independent of the channel noise distribution, and the channel coding theorem is independent of the source data distribution. Therefore, according to Shannon’s source-channel coding theorem, the conclusion is reached that if the original source is transmitted without distortion, separated source-channel coding (SSCC) is equivalent to JSCC, and both can achieve optimal transmission.
@@
However, these theorems are designed for lossless data transmission. When distortion is introduced to the original source, the communication process can no longer be adequately described by these three theorems alone. For semantic communication, semantics invariably depend on the existence of data, but data does not necessarily convey semantics. When the semantic encoder extracts semantic features from data, it removes data redundancies, making semantic communication potentially more efficient than classical communication methods. Building on this premise, we argues that semantic communication functions within the framework of rate distortion theory for data, achieving semantic fidelity in information transmission through the effective utilization of the knowledge base.
@@
In practical communication processes, source encoding is inherently lossy. This is due to the impracticality of accurately calculating the actual data entropy, especially for images. The challenge arises because the data being transmitted typically cannot effectively represent the entire population, and thus the estimated distribution does not accurately reflect the population distribution. Moreover, the goal of practical communication is generally to achieve better transmission performance in typical communication scenarios rather than to optimize over an infinitely long time. As a result, data is not considered from the perspective of a typical set, and the source does not exhibit the asymptotic equipartition property (AEP). The effectiveness of data compression and recovery is not simply measured by the bit error rate, but rather by evaluating the source encoding and decoding in terms of the content of the data within the current communication process. Therefore, even without considering channel matching, semantic has already been “applied” in practical communication processes.
III-B Joint and Separated Source-Channel Coding
Building on the investigation of the first question in section I, we further explores the second question. While the lossy transmission theory is not yet fully developed, existing proofs indicate that JSCC can still achieve optimality [44]. JSCC considers not only the source and channel noise distributions but also the relationship between source distortion and channel noise. In most cases of source distortion, the source encoding rate and channel encoding rate in SSCC lack closed-form analytical expressions, making it impossible to achieve optimality in practice [45]. Therefore, we propose a semantic communication model based on the JSCC framework that jointly optimizes knowledge base integration and effectiveness problem. This approach can provide advantages over SSCC when striving for optimal transmission under practical constraints.
@@ III-B Joint and Separated Source-Channel Coding
The data transmission described by the rate-distortion function no longer focuses solely on matching the source distribution $P(X)$ with the channel transition distribution $P(\hat{Y}|Y)$, as in CIT. Instead, it integrates the distortion function $d(x,\hat{x})$, source distribution $P(X)$, and channel transition distribution $P(\hat{Y}|Y)$. Consequently, optimal coding must not only minimize the distortion rate but also consider whether the channel noise can be leveraged as part of the distortion function. Subsequent studies [44] have elaborated on this approach.
@@
However, is JSCC truly indispensable for semantic communication? Unlike CIT, which benefits from the separation theorem, semantic communication lacks a well-established mathematical framework to support such endeavors. Recent studies [45] have shown that finding an analytical solution for the source distortion rate that perfectly matches a fixed channel coding rate, or vice versa, is unattainable. Nevertheless, if based on the JSCC framework, the IB theory [49] might provide the best interpretation for semantic communication systems. The IB describes the optimal representation of information during data compression and feature extraction. Its core idea is to retain the information most relevant to the target task (i.e., transmitting semantic information) while compressing redundant information from the raw data (i.e., reducing data transmission). Given input data $X$ and target task $\hat{X}$, we seek an intermediate representation $\hat{Y}$ that maximizes its predictive capability for $\hat{X}$ while minimizing its dependence on $X$. This can be achieved by minimizing Eq. 3:
$\mathcal{L}_{\text{IB}}[p(z|x)]=I(X;\hat{Y})-\beta I(\hat{Y};\hat{X}).$ (3)
@@
The IB problem actually corresponds to the semantic communication model as depicted in Fig.2. It is the proposed semantic communication model based on JSCC. While semantic information $S$ is transmitted in the data space, its theoretical measurement occurs in the semantic space. The data transmission model resembles the classical communication model, but with SSCC replaced by JSCC, potentially providing a better alignment between source data distortion $d(X;\hat{X})$ and channel noise. Additionally, semantic communication involves the deployment of a knowledge base and the consideration of communication objectives within the semantic space. The knowledge base, serving as a repository of prior information, bolsters the semantic encoding and decoding processes. Furthermore, for diverse communicative tasks, semantic information may also introduce distortions $d_{D}(S;\hat{S})$, retaining only the information relevant to the task, which suggests that semantic communication has the potential to further reduce data transmission. Consequently, the realization of semantic communication means we aim to ensure that the received information contains as much of the original semantic meaning as possible, which corresponds to the process of maximizing $\beta I(\hat{Y};\hat{X})$ in Eq.3. While achieving optimal semantic communication requires a delicate balance among various factors, including data distortions, source distribution, channel transition distribution, the knowledge base, and semantic distortions. It means we want the received information to be as compact as possible, which corresponds to the process of minimizing $I(X;\hat{Y})$ in Eq.3.
Traditional statistical theories lack the necessary tools to analyze such complex relationships, leveraging JSCC based on ANNs may offer a viable approximation. In deep learning, the Variational Information Bottleneck (VIB)  [50] is typically employed as a loss function to train ANNs and achieve IB-like results. The VIB loss function is given by:
$\mathcal{L}_{\text{VIB}}=$ $~{}\mathbb{E}_{x\sim p(x)}\mathbb{E}_{\hat{y}\sim p(\hat{y}|x)}\left[-\log q(\hat{x}|\hat{y})\right]$ (4)
$+\beta\cdot\mathbb{E}_{x\sim p(x)}\left[\text{KL}\big{(}p(\hat{y}|x)\parallel q(\hat{y})\big{)}\right].$
Notably, if we set $q(\hat{y})$ to be a standard normal distribution, and replace $\hat{y}$ with $z$, that is the standard latent variable symbol in VAEs, Eq. 4 becomes nearly identical to the loss function of a VAE, with the addition of a hyperparameter $\beta$. In the next section, we present our VSCC model derived through variational inference rather than from the IB perspective. This approach is taken because the IB theory primarily provides conceptual guidance rather than specific methodologies for modeling ANN loss functions. Variational inference offers a more effective way to incorporate channel modeling into the encoder and explain its role in communication.
III-C Variational Source-Channel Coding for Channel Matching
The exploration of the two preceding questions elucidates that effective semantic communication requires explicit channel modeling in the JSCC. Since semantic communication inherently involves data distortion to eliminate redundant information unnecessary for semantic expression, lossy encoding is applied at the source. According to the optimal theorem for JSCC under lossy conditions, the lossy source in semantic communication could be achieved through the channel noise. Consequently, the channel no longer acts as an obstacle that necessitates additional data redundancy, such as channel encoding, to offset its impact, but instead becomes an integral component of the semantic coding process. To effectively utilize channel for distorting source data, appropriate matching between the channel noise distribution and the source data distribution is necessary. This alignment essentially transforms the original source data distribution into a new distribution, facilitating controlled distortion of data through the noise distribution, while simultaneously preserving the essential characteristics of the original data distribution. Nevertheless, the new distribution may lack a closed-form analytical expression, because data distribution transformation has always been a challenge in statistics [46]. Fortunately, research on ANNs offers effective approximation methods for addressing these challenges [47][48]. Building on variational inference and VIB, this paper proposes a VSCC method that integrates channel state information into the training function of semantic communication, thereby achieving the necessary alignment process.